The arithmetic mean of two positive numbers a and b (a>b) is twice their geometric mean , then a:b is

Grade:12th pass

1 Answers

Nicho priyatham
8 years ago
  • from given we get
         a+b=4(ab)0.5  
  • squaring on both sides we get
            a2+b2-14ab=0
  • by dividing both sides by b2 we get (put m=a/b)
              m-14m+1=0
solve the equation to get value of a:b